OpenStudy (mergirlblue):

write an equation for the translation of (x-2)^2 + (y+1)^2 = 16 by four units left and 6 units up?

OpenStudy (studygurl14):

The equation of the circle has the form \(\Large (x-h)^2+(y-k)^2=r^2\) where h represents horizontal translation and k represents vertical translation. To translate \(\large (x-2)^2+(y+1)^2=16\) 4 units left and 6 units up, you need to subtract 4 from the h value and add 6 units to the k value.

OpenStudy (mergirlblue):

okay, let me try.. hold on

OpenStudy (mergirlblue):

what is r^2? 16^2 ?

OpenStudy (studygurl14):

igrnore that part. It's not important in this case

OpenStudy (mergirlblue):

i got (x-6)^2 + (y+7)^2 but that isn't one of my choices

OpenStudy (studygurl14):

What are your answer choices?

OpenStudy (mergirlblue):

(x+2)^2 + (y-5)^2, (x-2)^2 + (y-5)^2, (x+5)^2 + (y-2)^2, and (x+4)^2 + (y-5)^2

OpenStudy (studygurl14):

Hmm..odd. I don't know. @Michele_Laino

OpenStudy (mergirlblue):

neither do i

OpenStudy (michele_laino):

such traslation is: \[\Large \begin{gathered} x \to x + 4 = X \hfill \\ y \to y + 6 = Y \hfill \\ \end{gathered} \] where \(X,\;Y\) are the new variables

OpenStudy (mergirlblue):

I don't understand what you mean

OpenStudy (mergirlblue):

@michele_laino

OpenStudy (michele_laino):

please replace, into the original equation, \(x,y\) with \(X-4,Y-6\) respectively, what do you get?

OpenStudy (michele_laino):

for example, we can write this: \(x-2--->X-4-2=X-6\) and \(y+1---> Y-6+1=Y-5\)

OpenStudy (michele_laino):

hint: after that substitution, we get this equation: \[\Large {\left( {X - 6} \right)^2} + {\left( {Y - 5} \right)^2} = 16\] is it correct?

OpenStudy (studygurl14):

But wouldn't it be +6, not -6, since it is 6 unites up, not down

OpenStudy (michele_laino):

yes! And the new variable is \(Y=y+6\) where \(y\) is the old variable

OpenStudy (studygurl14):

1 + 6 = 7, so you'd get (y+7)^2, right? Except that not an answer choice